# Hatred Stems from Ignorance! Distillation of the Persuasion Modes in Countering Conversational Hate Speech

### v2

@@ 3.1 Counterspeech Baseline Datasets
Model Method Sexism Racism Religious bigotry Overall
@@
closed Llama-2 txt 0.96 0.96 0.96 0.96 0.95 0.95 0.90 0.88 0.88 0.93
@@
Open Llama-2 txt 0.70 0.51 0.30 0.78 0.76 0.77 0.70 0.61 0.53 0.57
@@
Table 2: Performance (binary F1) of training four models on two datasets for closed and open conversations: Using two settings—training on text only (txt) and text with the persuasion category (txt+pr)—we report the precision (P) and recall (R) along with F1 scores per topic and the average macro F1 overall score of the topics. The bold scores indicate the best macro F1 average for each model set (txt and text$+$pr). McNemar’s significant test illustrated between the results as ** (p $<$ .01), *** (p $<$ .001), and **** (p $<$ .0001).
Adhering to the conversational structure observed in multi-turn conversations, whose average length was seven or more turns, the final gold annotation count for single-turn conversations reached 1,500. The multi-turn conversation dataset has several topic categories. We merged the fine-grained topic labels to unify the hate speech labeling process in the closed conversation dataset. Meanwhile, the open (single-turn) conversation dataset lacked topic labels. Therefore, human annotators labeled the data by conversation topic and persuasion modes (for more details on the data preprocessing, Appendix A). Table  1 illustrates the distribution of the two datasets that we used for our experiments concerning hate speech and counterspeech, spanning the three topics.
Persuasion Mode Labeling
@@ 3.3 Counterspeech Generation
Figure 1: Proportion of persuasion modes (emotion, reason, and credibility) in human hate speech (HS) and counterspeech (CS), and in machine-generated counterspeech using GPT-3 and Llama 2
Figure 2: Proportion of persuasion modes in human counterspeech according to the topic
Figure 3: Proportion of persuasion modes in machine-generated counterspeech (GPT-3) according to the topic
Figure 4: Proportion of persuasion modes in machine-generated counterspeech (Llama 2) according to the topic
@@
Religious bigotry Closed(multi-turn) 1.08e-101 **** 3.55e-90 ****
Racism Closed(multi-turn) 1.22e-15 **** 1.04e-37 ****
Sexism Closed(multi-turn) 5.95e-113 **** 1.32e-124 ****
Religious bigotry Open(single-turn) 4.74e-22 **** 1.72e-22 ****
Sexism Open(single-turn) 4.81e-40 **** 1.88e-38 ****
Racism Open(single-turn) 1.08e-14 **** 0.007 ***
Table 3: Comparison of generative and human counterspeech. Using the chi-squared $X^{2}$ significant test, all comparisons are statistically significant (p $<$ 1.00), ** (p $<$ .01), *** (p $<$ .001), and **** (p $<$ .0001)
@@ 4 Results and Analysis
Figure 5: Multi-turn racism
Figure 6: Multi-turn sexism
Figure 7: Multi-turn religious bigotry
Entities identified by Riveter and power scores of personas across closed (multi-turn) conversations for three topics. A positive value indicates a stronger association with persuasion mode
Figure 8: Single-turn racism
Figure 9: Single-turn sexism
Figure 10: Single-turn religious bigotry
Entities identified by Riveter and power scores of personas across open (single-turn) conversations for three topics. A positive value indicates a stronger association with persuasion mode
@@
We analyzed the persuasion mode distribution by conversation interaction type and generative- and human-based counterspeech (see Figure 4). We also used the entity recognition and coreference resolution capabilities of Riveter  (Antoniak et al. 2023) to identify which characters are framed as having or lacking power and assess the social dynamics of personae in the conversations for each persuasion type (Figures 710). Overall, generative counterspeech persuades using emotions, while humans persuade using reason. As shown in Figure 4, for closed (multi-turn) interactions, the reason persuasion mode comprises approximately 45% of human counterspeech, compared to only 31% and 21% of generative counterspeech (GPT and Llama-2, respectively). Similarly, in open interactions, 51% of human counterspeech uses reason, compared with 33% and 19% in the GPT and Llama-2 generative models, respectively. Looking closely at the topic level, the reason persuasion mode is more often used in racism than in religious bigotry and sexism. Emotion, however, is more often used in counterspeech, especially in closed (multi-turn) interactions Figure 4. This finding can be further reiterated by reviewing the persona dynamics in the racism topic using Riveter (Figures 710). Using reasoning statements that contain ”immigration” and ”statistics” as entities have more powerful persona scores in conversations when discussing rationales to support counterpoints in closed (multi-turn) interactions (see Figure 7). These persona tend to be fact-based rationale (immigration rules and statistical studies), while persuasion in open interactions tends to use personal characteristics (e.g., skin color). Conversely, in conversations that used emotion-based persuasion, the pronoun ”you” had the highest persona dynamic power score. This mode in conversations tends to reflect shameful feelings shown by posters of hate speech. Interestingly, credibility has a slightly higher presence in human-authored counterspeech (8%) compared to machine-generated counterspeech (5%; Figure 4). As in closed (multi-turn) interactions, human-authored counterspeech supports the argument by using personal stories. For instance, on the topic of religious bigotry, conversations using the credibility persuasion mode have personal perspectives (with the first-person pronoun ”I”) when personal stories are recounted (Figure  7). On the topics of religious bigotry and sexism, conversations use credibility more than they do in the topic of racism, in both human-sourced and machine-generated counterspeech.
@@
The open interaction conversation data has the distinct feature of enabling the broader public (many users) to engage with the post compared to closed interaction conversations in which the dialog is exclusively between two users (such as on GPT and chat platforms). We use open interaction conversation data to examine the stance of the replies to counterspeech. In contrast to previous studies, such as  (Mathew et al. 2019), which considered the presence of replies and likes on a post as indicative of support to the counterspeech, our study implemented the nuance of assessing the element in counterspeech that makes it likely to receive support by focusing on the persuasion mode. Figure  11 illustrates the conversation flow between hate speech and counterspeech, as well as the stance of the replies received by the counterspeech. The general trend observed is that counterspeech receives replies that are against it (about 56%). Moreover, supportive replies tend to add context to further justify the counterpoint (22.5%). Most of the opposing replies (43%) do not directly attack the counterspeech; instead, they reject the counterpoint. We examined the distribution of supportive and opposing stances using the persuasion mode and present this in Figure 12. There is a noticeable tendency for replies to support counterspeech that uses the reason persuasion mode (50.4%), and replies against the counterspeech are at 74%. Replies that use the emotion persuasion mode and are opposed to the counterspeech are at 40.7%, and those that use the same persuasion mode and are supportive of the counterspeech are at 37.7%.
Figure 11: Flow of persuasion modes between hate speech (HS) and counternarratives (CNs) in single-turn conversations (open interactions)
Figure 12: Correlation between the persuasion modes and the reply’s stance
@@ 5 Discussion and Implication
Interaction types of the conversation clearly have nuanced effects on counterspeech’s exhibited persuasion mode. As shown in Figure 4, 51% of the counterspeech examples in open interactions use reason as the persuasion mode, while 45% of the counterspeech examples in closed interactions use reason. This may be partly due to the cost of public opinion and the users’ need to appear knowledgeable, which may drive their tendency to support their counterpoints with facts and reasoning. By studying the persona’s social dynamic in hate conversations using Riveter  (Antoniak et al. 2023), as shown in (Figures  7– 10), we can see this behavior, its high power score toward people of a targeted group (White or Black people), and its lower power score when using self-reference (I). This persuasion mode tends to make the the conversation focus on justifying counterpoints to the target group rather than expressing the reason behind the hate and linking it to the self as a valid reason for hate. We also show the effectiveness of using persuasion as a proxy to differentiate counterspeech from hate speech (Table 2), which highlights the persuasion mode’s ability to serve as an instrument of explanation for hate detection model interpretability. Overall, these results offer an overview of the counterspeech’s nuances and the conversation interaction type’s effect on how those confronting hate may incorporate persuasion modes to express their counterpoint and combat hate.
@@
Our results show a clear gap in the singular focus of current counterspeech works on emotion-based hate-speech analyses, which further confirms other works’ recent findings (Lahnala et al. 2022). For instance, the work by (Lahnala et al. 2022) highlights the need to look beyond emotion, as it shows that cognitive empathy mitigates hate better than emotional empathy. Figure  4 shows the clear tendency of human-written counterspeech to use reason in order to express a counterpoint (over 45%); this is comparable to the mere 19% of machine-generated counterspeech that uses reason. Further, emotion is the main persuasion mode. Generative counterspeech’s behavior highlights the vicious cycle of training hate speech detection models on emotion-based data and the complexity of incorporating reason as a factor in the current generated content. A few recent studies have highlighted machine-generated counterspeech’s limitations (e.g.,  (Mun et al. 2023; Sen et al. 2023)). Their conclusion focuses on a target-specific strategy and shows a lack of persuasive power in generated counterspeech when compared to human-written counterspeech. Our study’s findings intensify the call for studying counterspeech’s non-emotional factors as building blocks of hate mitigation.
@@
Open interaction conversations (such as those on social media) are a means to observe how the public perceives counterspeech from multiple users’ perspectives. Thus, we analyzed the replies to counterspeech and showed the stances of the replies toward each type of persuasion mode. The results in Figure 11 show that the conversational flow alternates between the counterspeech and the type of the reply’s stance. Approximately 59% of counterspeech examples received replies with opposing stances, especially if the examples used the emotional persuasion mode (representing about 40.7% of the opposing stances, as in Figure  12). A reply’s support stance is likely to favor the reason persuasion mode, as represented in about 50.4% of supportive replies. Since this study’s scope is to provide a high-level comparison of the conversation interaction types and persuasion mode, possible future research may incorporate a nuanced analysis of how users perceive counterspeech by a specific topic. Recent studies have investigated the role of conversation structure while studying hate detection. For instance, the work by  (Yi and Zubiaga 2023) surveyed session-based cyberbullying and discussed the structure of social media conversations to construct a cyberbullying dataset. Our study shows an empirical investigation that paves the way for future consideration of replies to counterspeech to enhance the interpretation of counterspeech effective factors. The size of the open interaction has its limitations in this study. Moreover, we justified the results using chi-squared $X^{2}$ significant test to validate the interpretation of the comparison between machine-generated and human-authored counterspeech (Table  3).
@@ Appendix A Preparation of the Baseline Datasets
Topic Unique dialoiuges our 50% set of dialogues
@@ Appendix B Annotation Setup
We use LabelBox as a platform to recruit and run annotation tasks. This platform enables conversation annotation, where the whole dialogue is shown to the annotator, with the possibility of labeling each conversation turn. We provide a thorough, detailed guideline for each interaction type (open/close) to the annotated, Figure 13, and Figure 14666link to a GitHub repository that contains the dataset and guideline of annotation will be provided. At this submission, we provide a link to the pdf file of the annotation guideline for the topic racism https://file.io/MLVfDZr7Ieii . A fair wage per hour has been provided to the team of two annotators as part of the Boost Tier payment setting of LabelBox.  [^7] https://labelbox.com/product/annotate/
@@ Appendix C Classification of the Hate Speech and Counterspeech Model Hyperparameter
Model Hyperparameter Customization
@@
Models P value
Closed Llama-2(txt) and Llama-2 (txt+pr) 2.02e-11 ****
BERT (txt) and BERT (txt+pr) 0.16
BiLSTM (txt) and BiLSTM (txt+pr) 3.64e-06 ****
SVM (txt) and SVM(txt+pr) 0.154
Open Llama-2(txt) and Llama-2 (txt+pr) 1.11e-16 ****
BERT (txt) and BERT (txt+pr) 0.735
BiLSTM (txt) and BiLSTM (txt+pr) 0.011 **
SVM (txt) and SVM(txt+pr) 0.00051 ***
Table 8: Significance of comparing the prediction pattern of the training model on the text as the baseline (txt) with the training model on text and persuasion (txt+pr) (All comparisons are statistically significant, alpha = 0.05.) (p $<$ 1.00), ** (p $<$ .01), *** (p $<$ .001), and **** (p $<$ .0001)
@@ Appendix D Generated Counterspeech Settings
Validation of the Weak Supervision Model (Benchmark Method)
Validation of the Weak Supervision Model (Human Validation)
To validate the result of weakly supervised labels for the generated counter, we selected 50 generated counters from each topic and the generative model and validated the persuasion label. An independent annotator (one of the authors) performed the validation process through the two datasets to ensure consistency in the label moderation process. As shown in Table  11, in the closed dataset, the overall error rate is relatively small: about 19% for GPT-3 and 12% for Llama 2. For the open (single-turn) dataset, the overall error rate is about 22% for GPT-3 and 16% for Llama 2.
Dataset Topic SEXISM RACISM RELIGIOUS BIGOTRY Overall
@@
Close and Open Llama-2-7B txt 0.20 0.23 0.13 0.31 0.23 0.22 0.20 0.25 0.19 0.20
@@
Table 9: Binary F1 score for training the models to classify the persuasion labels (emotion, reason, and credibility)
Dataset Models SEXISM RACISM RELIGIOUS BIGOTRY Overall
P R F1 P R F1 P R F1 F1
Closed (multi-turn) silver Llama-2-7B txt 0.94 0.94 0.94 0.93 0.92 0.92 0.92 0.92 0.92 0.93
@@
Table 10: Binary F1 score for training the models on the silver labels of closed (multi-turn) interactions
Dataset GModel Topic Error-persuasion
Closed(multi-turn) GPT3 Racism 22%
Sexism 24%
Religious bigotry 12%
Overall 19%
LLAMA2 Racism 2%
Sexism 22%
Religious bigotry 12%
Overall 12%
Open(single-turn) GPT3 Racism 24%
Sexism 22%
Religious bigotry 22%
Overall 22%
LLAMA2 Racism 12%
Sexism 20%
Religious bigotry 16%
Overall 16%
@@
Table 11: Error rate of weak persuasion labels in the generated counter by the two generative models GPT-3 and Llama 2 for the two datasets
Paper Checklist
  1. 1.
    For most authors…
    1. (a)
      Would answering this research question advance science without violating social contracts, such as violating privacy norms, perpetuating unfair profiling, exacerbating the socio-economic divide, or implying disrespect to societies or cultures? yes, we explain the motivation at § 1, and detailed the research implication at § 5
    2. (b)
      Do your main claims in the abstract and introduction accurately reflect the paper’s contributions and scope? yes, we detail the study contributions at § 1, and provide detailed discussion of the results for each research question at § 4
    3. (c)
      Do you clarify how the proposed methodological approach is appropriate for the claims made? yes, we start by providing the definition of counter speech and the hypothesis of persuasion mode at§ 1 and situate the contribution with the related work at § 2 RQ1 indicated at mainly at the methodology § 3.2, and at § 1. As for RQ2, we compare the different persuasion modes in humans and generated counters using the distribution of the persuasion in two types of data: closed (multi-turn) and open (single-turn) as explained at § 3.3, § 1 and § 4. For RQ3, we explain the flow of replies at § 3.1, and further details at § 5
    4. (d)
      Do you clarify what are possible artifacts in the data used, given population-specific distributions? yes, we report the detailed data selection and preparation at § 3.1 and Appendix A.
    5. (e)
      Did you describe the limitations of your work? yes, at discussion section§ 5
    6. (f)
      Did you discuss any potential negative societal impacts of your work? We detail the research implication at § 5 and ethical statement § 7
    7. (g)
      Did you discuss any potential misuse of your work? yes, discussion of research implication at § 5and ethical statement § 7
    8. (h)
      Did you describe steps taken to prevent or mitigate potential negative outcomes of the research, such as data and model documentation, data anonymization, responsible release, access control, and the reproducibility of findings? yes, at ethical statement § 7, and appendix A.
    9. (i)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? yes
  2. 2.
    Additionally, if your study involves hypotheses testing…
    1. (a)
      Did you clearly state the assumptions underlying all theoretical results? We provide empirical justification to the hypothesis and answer the research question.
    2. (b)
      Have you provided justifications for all theoretical results? We provide empirical justification to the hypothesis and answer the research question.
    3. (c)
      Did you discuss competing hypotheses or theories that might challenge or complement your theoretical results? We describe the main hypothesis at § 1 and situate the assumption with previous work § 2
    4. (d)
      Have you considered alternative mechanisms or explanations that might account for the same outcomes observed in your study? yes, full analysis of the results described at § 4, and we provide further details at appendix D for generated counter analysis.
    5. (e)
      Did you address potential biases or limitations in your theoretical framework? yes, we detail the discussion at § 5
    6. (f)
      Have you related your theoretical results to the existing literature in social science? yes, § 2, and research implication at § 5
    7. (g)
      Did you discuss the implications of your theoretical results for policy, practice, or further research in the social science domain? yes, we show in the discussion section a detailed implication of the empirical results § 5
  3. 3.
    Additionally, if you are including theoretical proofs…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? yes, we explain the hypothesis of the experiment design at § 1 and further justify it by sitting it with previous work § 2
    2. (b)
      Did you include complete proofs of all theoretical results? yes, full experimental results and discussion at sections § 4 and § 5
  4. 4.
    Additionally, if you ran machine learning experiments…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? detail explanation of the experiment hyperparameter or data annotation guideline are at Appendix A, B, C, and D.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? yes, we provide a detailed description of the training split at appendix C.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? yes, full detail of significant text and experiment design at appendix C.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? yes, at section § 3.2 and appendix C.
    5. (e)
      Do you justify how the proposed evaluation is sufficient and appropriate to the claims made? yes, we report the significant test of all of the main experiments, Macnamer for classification comparison at § 3.2, and appendix C. Also, for the counter generation, we provide detailed validation process at § 3.3, and appendix D.
    6. (f)
      Do you discuss what is “the cost“ of misclassification and fault (in)tolerance? yes, we report the significant test at appendix c, and indicates the limitation at § 5
  5. 5.
    Additionally, if you are using existing assets (e.g., code, data, models) or curating/releasing new assets, without compromising anonymity
    1. (a)
      If your work uses existing assets, did you cite the creators? yes, experiment set up at § 3.1
    2. (b)
      Did you mention the license of the assets? yes, models description at § 3.2, and appendix C. Also, dataset discussed at § 3.1
    3. (c)
      Did you include any new assets in the supplemental material or as a URL? yes, we include anonymous link to annotation guideline at appendix A
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? yes, we describe data labeling and collection at § 3.1 and appendix A
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? yes, as the domain of the study is counter hate speech analysis, thus we were cautious in including any official examples directly in the paper. instead, we use an optimal tool (Riveter) to analyze the discussion as we explained at ethical statement § 7
    6. (f)
      If you are curating or releasing new datasets, did you discuss how you intend to make your datasets FAIR ? yes,ethical statement § 7
    7. (g)
      If you are curating or releasing new datasets, did you create a Datasheet for the Dataset (see gebru2021datasheets)? yes, we prepare an electronic datasheet readme file at our private GitHub repository. We plan to share it after receiving the decision on the paper.
  6. 6.
    Additionally, if you used crowdsourcing or conducted research with human subjects, without compromising anonymity
    1. (a)
      Did you include the full text of instructions given to participants and screenshots? yes, detail annotation process at appendix A
    2. (b)
      Did you describe any potential participant risks, with mentions of Institutional Review Board (IRB) approvals? yes, detail annotation process at Appendix A, along with ethical statement § 7
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? yes, Annotation recruitment process at appendix A.
    4. (d)
      Did you discuss how data is stored, shared, and deidentified? yes, we discuss the data collection at appendix A. also we provide a description of the measures we used to protect the data at ethical statement § 7

### v3

@@ Abstract
00footnotetext: The title “Hatred Stems from Ignorance” reflects the study’s theoretical grounding in Aristotle’s Rhetoric, as interpreted by Cope, E. M., and Sandys, J. E. (2010). This foundation aligns with the insights of Ibn Rushd (Averroes), a renowned philosopher known for reintroducing Aristotle’s texts. The phrase is inspired by one of Ibn Rushd’s reflections on the roots of hatred and ignorance.
@@ 3.1 Counterspeech Baseline Datasets
Adhering to the conversational structure observed in multi-turn conversations, whose average length was seven or more turns, the final gold annotation count for single-turn conversations reached 1,500. The multi-turn conversation dataset has several topic categories. We merged the fine-grained topic labels to unify the hate speech labeling process in the closed conversation dataset. Meanwhile, the open (single-turn) conversation dataset lacked topic labels. Therefore, human annotators labeled the data by conversation topic and persuasion modes (for more details on the data preprocessing, Appendix A). Table  1 illustrates the distribution of the two datasets that we used for our experiments concerning hate speech and counterspeech, spanning the three topics  [^2] https://github.com/AbeerAldayel/counterspeech-persuasion-modes .
Persuasion Mode Labeling
Sexism Racism Religious Bigotry Macro Avg.
@@
Closed (multi-turn)
LLaMA-2 txt 0.96 0.96 0.96 0.96 0.95 0.95 0.90 0.88 0.88 0.93
@@
Open (single-turn)
LLaMA-2 txt 0.70 0.51 0.30 0.78 0.76 0.77 0.70 0.61 0.53 0.57
@@
[…]
@@ 3.3 Counterspeech Generation
(a) Overall persuasion mode distribution
(b) Human counterspeech by topic
(c) GPT-3 counterspeech by topic
(d) Llama 2 counterspeech by topic
Figure 1: Distribution of persuasion modes (in %) in counterspeech across overall and topic-specific contexts, comparing human, GPT-3, and LLaMA 2 responses in multi-turn (closed) and single-turn (open) interactions.
@@
Religious bigotry (mT) 1.08e-101 **** 3.55e-90 ****
Racism (mT) 1.22e-15 **** 1.04e-37 ****
Sexism (mT) 5.95e-113 **** 1.32e-124 ****
Religious bigotry (1T) 4.74e-22 **** 1.72e-22 ****
Sexism (1T) 4.81e-40 **** 1.88e-38 ****
Racism (1T) 1.08e-14 **** 0.007 ***
Table 3: Comparison of generative and human counterspeech for multi-turn (mT) and single-turn (1T). Using the chi-squared $X^{2}$ significant test, all comparisons are statistically significant (p $<$ 1.00), ** (p $<$ .01), *** (p $<$ .001), and **** (p $<$ .0001)
@@ 4 Results and Analysis
Multi-turn Single-turn
(a) Racism (Multi-turn)
(b) Racism (Single-turn)
(c) Sexism (Multi-turn)
(d) Sexism (Single-turn)
(e) Religious bigotry (Multi-turn)
(f) Religious bigotry (Single-turn)
Figure 2: Entities identified by Riveter and power scores of personas across multi-turn (mT) and single-turn (1T) conversations for three topics (racism, sexism, and religious bigotry). A positive value indicates a stronger association with the persuasion mode. Refer to caption Positive, Refer to caption Negative.
@@
We analyzed the persuasion mode distribution by conversation interaction type and generative- and human-based counterspeech (see Figure 1(a)). We also used the entity recognition and coreference resolution capabilities of Riveter  (Antoniak et al. 2023) to identify which characters are framed as having or lacking power and assess the social dynamics of personae in the conversations for each persuasion type (Figures 2(a)2(f)). Overall, generative counterspeech persuades using emotions, while humans persuade using reason. As shown in Figure 1(a), for closed (multi-turn) interactions, the reason persuasion mode comprises approximately 45% of human counterspeech, compared to only 31% and 21% of generative counterspeech (GPT and Llama-2, respectively). Similarly, in open interactions, 51% of human counterspeech uses reason, compared with 33% and 19% in the GPT and Llama-2 generative models, respectively. Looking closely at the topic level, the reason persuasion mode is more often used in racism than in religious bigotry and sexism. Emotion, however, is more often used in counterspeech, especially in closed (multi-turn) interactions Figure 1(b). This finding can be further reiterated by reviewing the persona dynamics in the racism topic using Riveter (Figures 2(a)2(b)). Using reasoning statements that contain ”immigration” and ”statistics” as entities have more powerful persona scores in conversations when discussing rationales to support counterpoints in closed (multi-turn) interactions (see Figure 2(a)). These persona tend to be fact-based rationale (immigration rules and statistical studies), while persuasion in open interactions tends to use personal characteristics (e.g., skin color). Conversely, in conversations that used emotion-based persuasion, the pronoun ”you” had the highest persona dynamic power score. This mode in conversations tends to reflect shameful feelings shown by posters of hate speech. Interestingly, credibility has a slightly higher presence in human-authored counterspeech (8%) compared to machine-generated counterspeech (5%; Figure 1(a)). As in closed (multi-turn) interactions, human-authored counterspeech supports the argument by using personal stories. For instance, on the topic of religious bigotry, conversations using the credibility persuasion mode have personal perspectives (with the first-person pronoun ”I”) when personal stories are recounted (Figure  2(e)). On the topics of religious bigotry and sexism, conversations use credibility more than they do in the topic of racism, in both human-sourced and machine-generated counterspeech.
Figure 3: Flow of persuasion modes between hate speech (HS) and counternarratives (CNs) in single-turn conversations (open interactions).
Figure 4: Correlation between the persuasion modes and the reply’s stance.
@@
The open interaction conversation data has the distinct feature of enabling the broader public (many users) to engage with the post compared to closed interaction conversations in which the dialog is exclusively between two users (such as on GPT and chat platforms). We use open interaction conversation data to examine the stance of the replies to counterspeech. In contrast to previous studies, such as  (Mathew et al. 2019), which considered the presence of replies and likes on a post as indicative of support to the counterspeech, our study implemented the nuance of assessing the element in counterspeech that makes it likely to receive support by focusing on the persuasion mode. Figure  4 illustrates the conversation flow between hate speech and counterspeech, as well as the stance of the replies received by the counterspeech. The general trend observed is that counterspeech receives replies that are against it (about 56%). Moreover, supportive replies tend to add context to further justify the counterpoint (22.5%). Most of the opposing replies (43%) do not directly attack the counterspeech; instead, they reject the counterpoint. We examined the distribution of supportive and opposing stances using the persuasion mode and present this in Figure 4. There is a noticeable tendency for replies to support counterspeech that uses the reason persuasion mode (50.4%), and replies against the counterspeech are at 74%. Replies that use the emotion persuasion mode and are opposed to the counterspeech are at 40.7%, and those that use the same persuasion mode and are supportive of the counterspeech are at 37.7%.
@@ 5 Discussion and Implication
Interaction types of the conversation clearly have nuanced effects on counterspeech’s exhibited persuasion mode. As shown in Figure 1(a), 51% of the counterspeech examples in open interactions use reason as the persuasion mode, while 45% of the counterspeech examples in closed interactions use reason. This may be partly due to the cost of public opinion and the users’ need to appear knowledgeable, which may drive their tendency to support their counterpoints with facts and reasoning. By studying the persona’s social dynamic in hate conversations using Riveter  (Antoniak et al. 2023), as shown in (Figures  2(a)– 2(f)), we can see this behavior, its high power score toward people of a targeted group (White or Black people), and its lower power score when using self-reference (I). This persuasion mode tends to make the the conversation focus on justifying counterpoints to the target group rather than expressing the reason behind the hate and linking it to the self as a valid reason for hate. We also show the effectiveness of using persuasion as a proxy to differentiate counterspeech from hate speech (Table 2), which highlights the persuasion mode’s ability to serve as an instrument of explanation for hate detection model interpretability. Overall, these results offer an overview of the counterspeech’s nuances and the conversation interaction type’s effect on how those confronting hate may incorporate persuasion modes to express their counterpoint and combat hate.
@@
Our results show a clear gap in the singular focus of current counterspeech works on emotion-based hate-speech analyses, which further confirms other works’ recent findings (Lahnala et al. 2022). For instance, the work by (Lahnala et al. 2022) highlights the need to look beyond emotion, as it shows that cognitive empathy mitigates hate better than emotional empathy. Figure  1(a) shows the clear tendency of human-written counterspeech to use reason in order to express a counterpoint (over 45%); this is comparable to the mere 19% of machine-generated counterspeech that uses reason. Further, emotion is the main persuasion mode. Generative counterspeech’s behavior highlights the vicious cycle of training hate speech detection models on emotion-based data and the complexity of incorporating reason as a factor in the current generated content. A few recent studies have highlighted machine-generated counterspeech’s limitations (e.g.,  (Mun et al. 2023; Sen et al. 2023)). Their conclusion focuses on a target-specific strategy and shows a lack of persuasive power in generated counterspeech when compared to human-written counterspeech. Our study’s findings intensify the call for studying counterspeech’s non-emotional factors as building blocks of hate mitigation.
@@
Open interaction conversations (such as those on social media) are a means to observe how the public perceives counterspeech from multiple users’ perspectives. Thus, we analyzed the replies to counterspeech and showed the stances of the replies toward each type of persuasion mode. The results in Figure 4 show that the conversational flow alternates between the counterspeech and the type of the reply’s stance. Approximately 59% of counterspeech examples received replies with opposing stances, especially if the examples used the emotional persuasion mode (representing about 40.7% of the opposing stances, as in Figure  4). A reply’s support stance is likely to favor the reason persuasion mode, as represented in about 50.4% of supportive replies. Since this study’s scope is to provide a high-level comparison of the conversation interaction types and persuasion mode, possible future research may incorporate a nuanced analysis of how users perceive counterspeech by a specific topic. Recent studies have investigated the role of conversation structure while studying hate detection. For instance, the work by  (Yi and Zubiaga 2023) surveyed session-based cyberbullying and discussed the structure of social media conversations to construct a cyberbullying dataset. Our study shows an empirical investigation that paves the way for future consideration of replies to counterspeech to enhance the interpretation of counterspeech effective factors. The size of the open interaction has its limitations in this study. Moreover, we justified the results using chi-squared $X^{2}$ significant test to validate the interpretation of the comparison between machine-generated and human-authored counterspeech (Table  3).
@@ Appendix A Preparation of the Baseline Datasets
Topic Unique dialogues 50% set of dialogues
@@ Appendix B Annotation Setup
We use LabelBox as a platform to recruit and run annotation tasks. This platform enables conversation annotation, where the whole dialogue is shown to the annotator, with the possibility of labeling each conversation turn. We provide a thorough, detailed guideline for each interaction type (open/close) to the annotators, Figure 5, and Figure 6777https://github.com/AbeerAldayel/counterspeech-persuasion-modes . A fair wage per hour has been provided to the team of two annotators as part of the Boost Tier payment setting of LabelBox.  [^8] https://labelbox.com/product/annotate/
@@ Appendix C Classification of the Hate Speech and Counterspeech Model Hyperparameter
Model Hyperparameter Value
@@
Dataset Model Comparison p-value
Closed (mT) LLaMA-2 (txt vs. txt+pr) 2.02e-11 ****
BERT (txt vs. txt+pr) 0.160
BiLSTM (txt vs. txt+pr) 3.64e-06 ****
SVM (txt vs. txt+pr) 0.154
Open (1T) LLaMA-2 (txt vs. txt+pr) 1.11e-16 ****
BERT (txt vs. txt+pr) 0.735
BiLSTM (txt vs. txt+pr) 0.011 **
SVM (txt vs. txt+pr) 0.00051 ***
[…]
@@ Appendix D Generated Counterspeech Settings
Dataset Model Input Sexism Racism Religious Overall
@@
Gold-Labeled (Closed + Open)
LLaMA-2-7B txt 0.20 0.23 0.13 0.31 0.23 0.22 0.20 0.25 0.19 0.20
@@
Silver-Labeled (Closed - Multi-turn)
LLaMA-2-7B txt 0.94 0.94 0.94 0.93 0.92 0.92 0.92 0.92 0.92 0.93
@@
Table 9: Precision (P), Recall (R), and F1 scores for persuasion mode classification across datasets. Top: models trained on gold-labeled data (closed + open). Bottom: models trained on silver-labeled data (multi-turn only).
Validation of the Weak Supervision Model (Benchmark Method)
Validation of the Weak Supervision Model (Human Validation)
To validate the result of weakly supervised labels for the generated counter, we selected 50 generated counters from each topic and the generative model and validated the persuasion label. An independent annotator (one of the authors) performed the validation process through the two datasets to ensure consistency in the label moderation process. As shown in Table  10, in the closed dataset, the overall error rate is relatively small: about 19% for GPT-3 and 12% for Llama 2. For the open (single-turn) dataset, the overall error rate is about 22% for GPT-3 and 16% for Llama 2.
Model Racism Sexism Religious Overall
Closed ( Multi-turn)
GPT-3 22% 24% 12% 19%
LLaMA 2 2% 22% 12% 12%
Open ( Single-turn)
GPT-3 24% 22% 22% 22%
LLaMA 2 12% 20% 16% 16%
@@
Table 10: Error rates (%) in weak persuasion labels for machine-generated counterspeech across models.